OpenStudy (anonymous):

can someone help me solve sq rt 4 of 64^7 divided by sq rt 4 of x

OpenStudy (anonymous):

can u type the eqn out? lol, sorry

OpenStudy (anonymous):

\[\sqrt[4]{64x^7}\times \sqrt[4]{x}\]

OpenStudy (anonymous):

you can combine functions\[\sqrt[4]{64x^7}\times \sqrt[4]{x}\] =\[4\times x ^{7/4} \times x ^{1/4}\] \[4x ^{8/4}\] because of the rule of powers\[=4x^2\] done!